a father is 42 years old. His children were 6 years, 8 years and 12 years. How old will the father when his age will be equal to the sum of the ages of his three children?
Answer by ankor@dixie-net.com(22740)

a father is 42 years old. His children were 6 years, 8 years and 12 years.
How old will the father be when his age will be equal to the sum of the ages of his three children?
:
Let y = no. of years for this to be true
:
(y+6) + (y+8) + (y+12) = y + 42
3y + 26 = y + 42
3y - y = 42 - 26
2y = 16
y = 8
:
He will 42 + 8 = 50 yrs old
:
You can check this, add 8 to each child, see if it adds up to 50